# Towards Realistic Evaluation of Commit Message Generation by Matching Online and Offline Settings

### v1

@@ Abstract
Our results indicate that edit distance exhibits the highest correlation, whereas commonly used similarity metrics such as BLEU and METEOR demonstrate low correlation. This contradicts the previous studies on similarity metrics for CMG, suggesting that user interactions with a CMG system in real-world settings differ significantly from the responses by human labelers operating within controlled research environments. While our findings are tied to the CMG model we used, and the results may vary for models with cardinally different outputs, our proposed framework is relatively lightweight in terms of required human effort, and we release all the code and the dataset to support future research in the field of CMG: https://jb.gg/cmg-evaluation.
@@ I Introduction
Building off this idea, we propose to employ multiple pairs of generated messages $G_{i}$ and their edited versions $E_{i}$ for each commit $C$ in the offline metric selection process. Each related pair, where $E_{i}$ was obtained directly from $G_{i}$, provides us with an intuitive and practical metric—the effort that the editing process took—thus providing the online evaluation of the $G_{i}$. At the same time, edited messages $E_{j\neq i}$ obtained from other generated messages for the same commit and the original commit message $O$ scraped from the VCS provide reference messages for $G_{i}$, suitable for offline evaluation. Hence, we can evaluate each generated message $G_{i}$ both online (against $E_{i}$) and offline (against $E_{j\neq i}$ and $O$). This property allows us to gauge how good different metrics for the offline evaluation will estimate the online performance of the model.
@@
As part of this effort, we developed an app for collecting the expert markup for the generated commit messages and used it to extend a CMG dataset from Long Code Arena [11] with multiple $(G,E)$ pairs per commit. We generated 15 commit messages with GPT-4 [12], collected 57 human edits for those generated messages, and extended the dataset by synthetically generating 599 more $(G,E)$ pairs. We assess the distribution of our dataset in relation to logs from real users of a CMG system within PyCharm, a Python IDE from JetBrains, and demonstrate that our dataset remains indicative of real user behavior.
@@
  • We describe the metric selection method and share our findings (Section IV).
@@ II Framework
Most existing datasets for offline evaluation of CMG models are collected from open sources, e.g., GitHub. There are two key entities associated with each commit in the VCS: the set of code changes $C$ and an original commit message $O$ written by developers (the upper part of Figure 1). The quality of the CMG model is then obtained by calculating $\mathbf{m(G,O)}$, where $\mathbf{m}$ is a text similarity metric, and $G$ is a model-generated commit message. However, during real usage of a CMG system, users often edit the generated message $G$ before submitting it to VCS, thus introducing the third entity for a commit—the edited message $E$. We illustrate this idea with motivating examples: for instance, a user may want to shorten the generated message (in Figure 1, $E_{1}$ and $E_{2}$ are shortened versions of $G_{1}$ and $G_{2}$, respectively) or to rewrite it to follow the conventions from their project (in Figure 1, $E_{3}$ is a version of $G_{3}$ adapted according to Conventional Commits [13] specification). Among the online metrics, a common choice is to estimate the effort the user edit took (e.g., time, edit distance [9]). Guided by this intuition, we propose to consider the correlation of $\mathbf{m(G,O)}$ with the estimation of the editing effort as the indication of how well metric $\mathbf{m}$ approximates the way real users perceive quality. Specifically, at JetBrains, $\mathbf{ED(G,E)}$—edit distance [9] between the generated messages $G$ and their user-edited versions $E$—is one of the important indicators of online quality of our CMG system. Hence, in our experiments, we use $\mathbf{ED(G,E)}$ as the online measure of CMG quality.
@@ III Dataset
The process of dataset collection is illustrated in Figure 2. We implemented a web application for labeling (Section III-A) and collecting message edits from human experts (Section III-B, illustrated by Figure 2(a)). Given the manual effort of rewriting commit messages and the high quality of synthetic data generation reported in numerous research works [14], we then propose two different ways to synthetically augment the dataset, backward and forward (Section III-C, illustrated by Figure 2(b) and Figure 2(c), respectively). Finally, we show that our dataset is indicative of real user behavior by comparing the distribution of edits with user logs from PyCharm, a Python IDE developed by JetBrains (Section III-D).
@@ III-A Dataset Collection Application
Figure 3: A screenshot of our web application for commit message edits collection. On the left, the assessors are presented with a set of code changes from the current commit in a diff format. On the right, there is a window with a model-generated commit message to be edited; all changes in this window are tracked. Additionally, there is a Help toggle with a labeling instruction on top and a Commit summary toggle with extra information about the current project, aimed to provide more context for the code changes. The web application is publicly available: https://huggingface.co/spaces/JetBrains-Research/commit-message-editing
@@
So, to collect a dataset closely simulating the intended user behavior, we prepare and release to the open-source a web application [^1] Commit Message Editing web application: https://huggingface.co/spaces/JetBrains-Research/commit-message-editing implemented using the Gradio framework [15]. The graphical interface of the application is presented in Figure 3. The application provides an assessor with the diff—the set of code changes from the commit—and a window with the model-generated message $G$ to be edited. The application collects all the keystrokes made in the editing window, including content, location and timestamp for each change; therefore, the user typing can be fully reproduced from the collected data.
@@ III-B Dataset Collection
The labeling was carried out by eight researchers from JetBrains Research and three PhD students with several years of programming experience in the industry. All participants were given the same set of commits, but the commits were randomly shuffled for each participant to avoid possible bias from the predefined order.
As the source of the commits, we selected the Commit Message Generation benchmark from Long Code Arena [11]. The benchmark features $163$ commits from Python repositories, all of which are manually confirmed to have complex and meaningful code changes akin to those developers encounter in their daily work. For model-generated commit messages $G$ that experts need to edit, we employed the GPT-4 results released by the authors of the benchmark [16]. We randomly selected a single set of $15$ commits that we present to each participant to achieve a higher number of labels per sample.
We argue that having multiple labels per sample offers significant advantages. For each model-generated message $G$, all edited messages except for the related can serve as high-quality references. Drawing inspiration from classical metrics used in reference-based evaluation, such as BLEU [17] or ROUGE [18], we believe that employing multiple references can lead to more stable and reliable evaluations. Moreover, we intentionally presented the labelers with an abstract instruction of “edit the commit message so that it is good enough to publish to VCS” to collect different types of edits, so having only one edited sample per commit would have inadvertently introduced a bias.
In total, we collected $57$ manually edited samples, with $3.8$ samples per commit on average. Our expert-labeled dataset is available online. [^2] Expert-labeled dataset: https://huggingface.co/datasets/JetBrains-Research/commit-msg-edits
@@ III-C Dataset Extension
Manually editing commit messages takes significant effort. Since we intend our framework to be useful for practitioners evaluating CMG models, it is important to maintain resource efficiency. With that in mind, we explore the possibility of synthetically extending our dataset  [14]. To ensure this approach does not introduce bias, we compare the distribution of the synthetic subset with real interaction logs from PyCharm, a Python IDE developed by JetBrains, as detailed in Section III-D.
To imitate human actions, we employ a powerful LLM GPT-4 Turbo [19] together with the in-context learning (ICL) technique. Specifically, we propose two different approaches to extend our dataset. They are illustrated in Figures 2(b) and 2(c) and detailed below.
For backward generation (Figure 2(b)), LLM is expected to “invert” the edit performed by a human expert. In other words, the LLM is given the manually edited version of a commit message $E$, and its task is to “restore” the original model-generated message $G$. To aid generation, we emply in-context learning and supply the model with several examples of pairs $E,G$ for different commits. We denote the result of such backward generation $G^{\prime}$—a synthetic model-generated commit message. To eliminate the samples that are too far from the edit distributions from the expert-labeled dataset, we discard samples that have more than $T_{b}\%$ of content added during this step, where $T_{b}$ is a hyperparameter to be set.
For forward generation (Figure 2(c)), LLM imitates precisely what human experts did during labeling, i.e., rewrites a model-generated commit message (both usual $G$ and synthetic $G^{\prime}$) to $E$. Here, as ICL examples of such conversion, we again used pairs of $G,E$ for different commits. The result is denoted $E^{\prime}$—a synthetic edited commit message. To alleviate the samples that are too far from the edit distributions from the expert-labeled dataset, we discard samples that have more than $T_{f}\%$ of content removed during this step, where $T_{f}$ is a hyperparameter to be set.
@@
Figure 4: Distribution of the $\mathbf{ED(G,E)}$ values for the different subsets of our dataset and for the PyCharm users logs over one month in April-May 2024, where $\mathbf{ED(G,E)}$ is the edit distance between model-generated messages $\mathbf{G}$ and their edited versions $\mathbf{E}$. Note that we scale the PyCharm logs $\mathbf{ED}$ values to adjust for the differences in messages’ lengths and discard the samples with edit distance $\mathbf{=0}$.
The detailed statistics for the resulting dataset are provided in Table I. Our synthetic extension of the dataset is also available online. [^3] Synthetic dataset, featuring backward-generated and forward-generated samples: https://huggingface.co/datasets/JetBrains-Research/synthetic-commit-msg-edits
@@ III-D Data Validation
As part of the Early Access Program for IntelliJ-based IDEs, JetBrains collects anonymous usage statistics. This includes metrics like $\mathbf{ED(G,E)}$ — the edit distance between the commit message $G$ generated by the AI Assistant plugin and the message $E$ that actually gets stored in the VCS. To ensure that our expert-labeled dataset and its synthetic extension remain representative of real usage, we compare the distributions of $\mathbf{ED(G,E)}$ in our dataset to the same metrics computed for actual users of PyCharm, a Python IDE developed by JetBrains.
Figure 4 shows the distribution of edit distance values for the PyCharm users over one month in April-May 2024, as well as for the different subsets of our dataset.
We note that the majority of the samples in PyCharm user logs have very low edit distance values ($78\%$ of all samples have an edit distance value equal to 00), while in our dataset, there are no such samples at all. One possible reason is that in such cases, the quality of the generated commit messages is sufficient for the users to keep them as is. On the other hand, commit messages might not be as important compared to the effort required to verify and edit them. Though there are a lot of samples with edit distance equal to 00, they come from only $33$% of users. As we do not want to consider the cases with low user engagement, we remove the samples with a zero edit distance, leaving us with $22$% of total samples coming from $67$% of total users.
@@
As observed in Figure 4, real PyCharm users tend to perform less editing than experts in our dataset. However, we note that the peak of the edit distance distribution, $\approx 400$, for the user logs data is relatively close to the peak for the golden dataset. Therefore, the data we collected is, to a certain extent, representative of real-life user behavior.
@@ IV Metrics Study
Figure 5: Overview of the relations between different $(G,E)$ pairs in our dataset. The pairs connected with the lines are the related pairs. Any pair $(G,E)$ for a given commit message $C$ that is not related is conditionally independent. We employ related pairs to calculate online metrics and conditionally independent pairs to calculate offline metrics during the metric selection process.
@@ IV-A Motivation
This way, for each commit, we obtain both $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$, and $m(G_{i},E_{j}),E_{j}\in\text{conditionally independent}(G_{i})$, where $m$ is a text similarity metric and $ED$ is the edit distance metrics that we employ as an estimation of the user effort the edit took, as it is one of the important online metrics computed for the actual CMG system at JetBrains. We do not use the time required to perform an edit because it is available only for the expert-labeled part of our dataset and not for the synthetic.
The $ED$ values better describe the quality of the CMG system from the point of view of its end users. However, obtaining them involves manually editing model-generated messages, so the process is infeasible to scale to large datasets or repeat numerous times, in contrast with reference-based evaluation with automatic metrics. Instead, we propose to use these values to select the most suitable metrics $m$ for a reference-based evaluation. Put another way, our core assumption is that the correlation of $m$ with $ED$ is indicative of how good $m$ is. Additionally, following the design of many classical metrics, e.g., BLEU [17] or ROUGE [18], we employ the presence of multiple references in our dataset and aggregate $m$ values. We argue that the conditionally independent messages $E$ might serve as references of high quality for each message $G$ as compared to the original commit message $O$, considering issues with the quality of commit messages coming from the open sources [1].
@@ IV-B Method
On the other hand, we compute $m$ for all conditionally independent pairs $(G_{i},E_{j})$, where $E_{j}\in\text{conditionally independent}(G_{i})$, i.e., it is an edited version of some other message $G_{j}$ and is not directly related to $G_{i}$. As an aggregation function $f_{agg}$, similarly to BLEU and ROUGE, we use the maximum if the higher values of $m$ indicate better quality or the minimum if the lower values of $m$ indicate better quality, e.g., for edit distance.
Finally, we compute the correlation between these two calculated values:

### v2

@@ Abstract
Our results indicate that edit distance exhibits the highest correlation with the online metric, whereas commonly used similarity metrics such as BLEU and METEOR demonstrate low correlation. This contradicts the previous studies on similarity metrics for CMG, suggesting that user interactions with a CMG system in real-world settings differ significantly from the responses by human labelers within controlled environments. While our findings are tied to the CMG model we used, and the results may vary for models with cardinally different outputs, our proposed framework is relatively lightweight in terms of required human effort, and we release all the code and the dataset to support future research in the field: https://jb.gg/cmg-evaluation.
@@ I Introduction
Building off this idea, we propose to employ multiple pairs of generated messages $G_{i}$ and their edited versions $E_{i}$ for each commit $C$ in the offline metric selection process. Each related pair, where $E_{i}$ was obtained directly from $G_{i}$, provides us with an intuitive and practical metric—the effort that the editing process took—thus providing the online evaluation of the $G_{i}$. At the same time, edited messages $E_{j\neq i}$ obtained from other generated messages for the same commit, as well as the original commit message $O$ scraped from the VCS, provide reference messages for $G_{i}$, suitable for offline evaluation. Hence, we can evaluate each generated message $G_{i}$ both online (against $E_{i}$) and offline (against $E_{j\neq i}$ and $O$). This allows us to gauge how good different metrics for the offline evaluation will estimate the online performance of the model.
@@
As part of this effort, we developed an app for collecting the expert markup for the generated commit messages and used it to extend a CMG dataset from Long Code Arena [11] with multiple $(G,E)$ pairs per commit. For 15 commit messages generated by GPT-4 [12], we collected 57 edits from human experts and extended the dataset by synthetically generating 599 more $(G,E)$ pairs. We assess the distribution of our dataset in relation to logs from real users of a CMG system within PyCharm, a Python IDE from JetBrains, and demonstrate that our dataset remains indicative of real user behavior.
@@
  • We describe the method for metric selection and share our findings (Section IV).
@@ II Framework
Most existing datasets for offline evaluation of CMG models are collected from open sources, e.g., GitHub. There are two key entities associated with each commit in the VCS: the set of code changes $C$ and an original commit message $O$ written by developers (the upper part of Figure 1). The quality of the CMG model is then obtained by calculating $\mathbf{m(G,O)}$, where $\mathbf{m}$ is a text similarity metric, and $G$ is a model-generated commit message. However, during real usage of a CMG system, users often edit the generated message $G$ before submitting it to VCS, thus introducing the third entity for a commit—the edited message $E$. We illustrate this idea with motivating examples: for instance, a user may want to shorten the generated message (in Figure 1, $E_{1}$ and $E_{2}$ are shortened versions of $G_{1}$ and $G_{2}$, respectively) or to rewrite it to follow the conventions from their project (in Figure 1, $E_{3}$ is a version of $G_{3}$ adapted to the Conventional Commits [13] specification). Among the online metrics, a common choice is to estimate the effort the user edit took (e.g., time, edit distance [9]). Guided by this intuition, we propose to consider the correlation of $\mathbf{m(G,O)}$ with the estimation of the editing effort as the indication of how well metric $\mathbf{m}$ approximates the way real users perceive quality. Specifically, at JetBrains, $\mathbf{ED(G,E)}$—edit distance [9] between the generated messages $G$ and their user-edited versions $E$—is one of the important indicators of online quality of our CMG system. Hence, in our experiments, we use $\mathbf{ED(G,E)}$ as the online measure.
@@ III Dataset
The process of dataset collection is illustrated in Figure 2. We implemented a web application (Section III-A) for labeling and collecting message edits from human experts (Section III-B, illustrated by Figure 2a). Given the manual effort of rewriting commit messages and the high quality of synthetic data generation reported in numerous research works [14], we then propose two different ways to synthetically augment the dataset, backward and forward (Section III-C, illustrated by Figure 2b and Figure 2c, respectively). Finally, we show that our dataset is indicative of real user behavior by comparing the distribution of edits with user logs from PyCharm, a Python IDE developed by JetBrains (Section III-D).
@@ III-A Dataset Collection Application
Figure 3: A screenshot of our web application for the collection of commit message edits. On the left, the assessors are presented with a set of code changes from the current commit in a diff format. On the right, there is a window with a model-generated commit message to be edited, all changes in this window are tracked. Additionally, there is a Help toggle with a labeling instruction on top and a Commit summary toggle with extra information about the current project, aimed to provide more context for the code changes. The web application is publicly available [15].
@@
So, to collect a dataset closely simulating the intended user behavior, we prepare and release to the open-source a web application [15] implemented using the Gradio framework [16]. The graphical interface of the application is presented in Figure 3. The application provides an assessor with the diff—the set of code changes from the commit—and a window with the model-generated message $G$ to be edited. The application collects all the keystrokes in the editing window, including content, location and timestamp for each change, so the user typing can be fully reproduced from the collected data.
@@ III-B Dataset Collection
The labeling was carried out by eight researchers from JetBrains Research and three PhD students with several years of experience in the industry. All participants were given the same set of commits, but the commits were randomly shuffled to avoid possible bias from the predefined order.
As the source of the commits, we selected the Commit Message Generation benchmark from Long Code Arena [11]. It features $163$ commits from Python repositories, all of which are manually confirmed to have complex and meaningful code changes akin to those developers encounter in their daily work. For model-generated commit messages $G$ that experts need to edit, we employed the results from GPT-4 (gpt4-06-13), a powerful proprietary LLM from OpenAI [12], released by the authors of the benchmark [17]. The prompt and the code used to obtain the GPT-4-generated messages are available in the replication package of Long Code Arena [18].
We randomly selected a single set of $15$ commits that we present to each participant to achieve a higher number of labels per sample. We argue that having multiple labels per sample offers significant advantages. For each model-generated message $G$, all edited messages except for the related one can serve as high-quality references. Drawing inspiration from classical metrics used in reference-based evaluation, such as BLEU [19] or ROUGE [20], we believe that employing multiple references can lead to more stable and reliable evaluations. Moreover, we intentionally presented the labelers with an abstract instruction of “edit the commit message so that it is good enough to publish to VCS” to collect different types of edits, so having only one edited sample per commit would have inadvertently introduced a bias.
In total, we collected $57$ manually edited samples, with $3.8$ samples per commit on average. Our expert-labeled dataset is available online [21].
TABLE I: Summary of the collected dataset.
@@ III-C Dataset Extension
Manually editing commit messages takes significant effort. Since we intend our framework to be useful for practitioners, it is important to maintain resource efficiency. With that in mind, we explore the possibility of synthetically extending our dataset [14]. To ensure this approach does not introduce bias, we compare the distribution of the synthetic subset with real interaction logs from PyCharm, a Python IDE developed by JetBrains, as detailed in Section III-D.
To imitate human actions, we employ GPT-4 Turbo [22] (gpt-4-1106-preview), a powerful LLM from OpenAI, together with the in-context learning (ICL) technique. Specifically, we propose two different approaches to extend our dataset, illustrated in Figures 2b and 2c and detailed below. We release the code used to obtain the synthetic data [23].
For backward generation (Figure 2b), LLM is expected to “invert” the edit performed by a human expert. In other words, the LLM is given the manually edited version of a commit message $E$, and its task is to “restore” the original model-generated message $G$. To aid generation, we employ in-context learning and supply the model with several examples of pairs $E,G$ for different commits. We denote the result of such backward generation $G^{\prime}$—a synthetic model-generated commit message. To eliminate the samples that are too far from the edit distributions from the expert-labeled dataset, we discard samples that have more than $T_{b}\%$ of content added during this step, where $T_{b}$ is a hyperparameter to be set.
For forward generation (Figure 2c), LLM imitates precisely what human experts did during labeling, i.e., rewrites a model-generated commit message (both usual $G$ and synthetic $G^{\prime}$) to $E$. Here, as ICL examples of such conversion, we again used pairs of $G,E$ for different commits. The result is denoted $E^{\prime}$—a synthetic edited commit message. To alleviate the samples that are too far from the edit distributions from the expert-labeled dataset, we discard samples that have more than $T_{f}\%$ of content removed during this step, where $T_{f}$ is a hyperparameter to be set.
@@
Figure 4: Distribution of the $\mathbf{ED(G,E)}$ values for the different subsets of our dataset and for the PyCharm users logs over one month in April-May 2024, where $\mathbf{ED(G,E)}$ is the edit distance between model-generated messages $\mathbf{G}$ and their edited versions $\mathbf{E}$. Note that we scale the PyCharm logs $\mathbf{ED}$ values to adjust for the differences in messages’ lengths and discard the samples with edit distance equal to 0.
The detailed statistics for the resulting dataset are provided in Table I. Our synthetic extension of the dataset is also available online [24].
@@ III-D Data Validation
As part of the Early Access Program for IntelliJ-based IDEs, JetBrains collects anonymous usage statistics. This includes $\mathbf{ED(G,E)}$ — the edit distance between the commit message $G$ generated by the AI Assistant plugin and the message $E$ that actually gets stored in the VCS. To ensure that our expert-labeled dataset and its synthetic extension remain representative of real usage, we compare the distributions of $\mathbf{ED(G,E)}$ in our dataset to the same metrics computed for actual users of PyCharm, a Python IDE developed by JetBrains.
Figure 4 shows the distribution of edit distance values for the PyCharm users over one month in April-May 2024, as well as for the different subsets of our dataset. We note that the majority of the samples in PyCharm user logs have very low edit distance values ($78\%$ of all samples have an edit distance value equal to 00), while in our dataset, there are no such samples at all. One possible reason is that in such cases, the quality of the generated commit messages is sufficient for the users to keep them as is. On the other hand, commit messages might not be as important compared to the effort required to verify and edit them. Though there are a lot of samples with edit distance equal to 00, they come from only $33$% of users. As we do not want to consider the cases with low user engagement, we remove the samples with a zero edit distance, leaving us with $22$% of total samples coming from $67$% of total users.
@@
As observed in Figure 4, real PyCharm users tend to perform less editing than experts in our dataset. However, we note that the peak of the edit distance distribution, $\approx 400$, for the user logs data is relatively close to the peak for the golden expert-labeled dataset. Therefore, the data we collected is, to a certain extent, representative of real-life user behavior.
@@ IV Metrics Study
Figure 5: Overview of the relations between different $(G,E)$ pairs in our dataset. The pairs connected with the lines are the related pairs. Any pair $(G,E)$ for a given commit message $C$ that is not related is conditionally independent. We employ related pairs to calculate online metrics and conditionally independent pairs to calculate offline metrics.
@@ IV-A Motivation
This way, for each commit, we obtain both $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$, and $m(G_{i},E_{j}),E_{j}\in\text{conditionally independent}(G_{i})$, where $m$ is a text similarity metric and $ED$ is the edit distance metric that we employ as an estimation of the user effort the edit took, as it is one of the important online metrics computed for the actual CMG system at JetBrains. We do not use the time required to perform an edit because it is available only for the expert-labeled part of our dataset and not for the synthetic.
The $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$ values better describe the quality of the CMG system from the point of view of its end users. However, obtaining them involves manually editing model-generated messages, so the process is infeasible to scale to large datasets or repeat numerous times, in contrast with reference-based evaluation with automatic metrics. Instead, we propose to use these values to select the most suitable metrics $m$ for a reference-based evaluation. Put another way, our core assumption is that the correlation of $m$ with $ED(G_{i},E_{j}),E_{j}\in\text{related}(G_{i})$ is indicative of how good $m$ is. Additionally, following the design of many classical metrics, e.g., BLEU [19] or ROUGE [20], we use the multiple references in our dataset and aggregate $m$ values. We argue that the conditionally independent messages $E$ might serve as references of high quality for each message $G$ as compared to the original commit message $O$, considering issues with the quality of commit messages coming from the open sources [1].
@@ IV-B Method
On the other hand, we compute $m$ for all conditionally independent pairs $(G_{i},E_{j})$, where $E_{j}\in\text{conditionally independent}(G_{i})$, i.e., it is an edited version of some other message $G_{j}$ and is not directly related to $G_{i}$. As an aggregation function $f_{agg}$, similarly to BLEU and ROUGE, we use the maximum if the higher values of $m$ indicate better quality or the minimum if the lower values of $m$ indicate better quality, e.g., for edit distance. Finally, we compute the correlation between these two calculated values:
@@
Lastly, the primary online metric we consider is the edit distance between generated messages and their edited counterparts. However, we acknowledge that other metrics could be a valid choice for online evaluation as well. Our dataset and framework allows to easily replace $ED$ in $ED(G_{i},E_{j})$ where $E_{j}\in\text{related}(G_{i})$ with any other metric $m^{*}$. In addition to our primary experiments with $Q(m)$, we also consider the correlation coefficients:
$\begin{split}Q^{*}(m,{\color[rgb]{1,0,0}m^{*}}):=\text{corr }\Big{(}\text{mean}\big{(}\{{\color[rgb]{1,0,0}m^{*}}(G_{i},E_{j})\}_{E_{j}\in\text{related}(G_{i})}\big{)},\\ f_{agg}\big{(}\{m(G_{i},E_{j})\}_{E_{j}\in{\text{conditionally independent}(G_{i})}}\big{)}\Big{)}\end{split}$
We compute and release $Q^{*}(m,m^{*})$ for the set of 9 text similarity metrics listed above used both as an offline metric $m$ and as an online metric $m^{*}$ [23].